OpenStudy (anonymous):

Choose the correct simplification of the expression (8x2)3. 24x6 512x6 512x5 24x5

OpenStudy (anonymous):

24 x 6 , 8*3=24 2*3 = 6 24*6.

OpenStudy (anonymous):

i though it was 24x^5?

OpenStudy (agent47):

8=2^3 8*2=2^4 (2^4)^3=2^12

OpenStudy (agent47):

=4096

OpenStudy (agent47):

waittt what do you mean by x? Your whole power and multiplication notations are confusing

OpenStudy (anonymous):

dont i have to add the 3 and 2?

OpenStudy (agent47):

(8^2)^3? no u have to multiply them

OpenStudy (anonymous):

I thought x was multiply so my answer is wrong,

OpenStudy (anonymous):

you multiply the 8 by the 2 so its 8*8 and then add the 3 and 2. is that corret?

jimthompson5910 (jim_thompson5910):

(8x^2)^3 (8x^2)(8x^2)(8x^2) (8*8*8)(x^2*x^2*x^2) 512x^(2+2+2) 512x^6 So (8x^2)^3 simplifies to 512x^6

OpenStudy (anonymous):

Is the original expression: \[(8^{2})^{3}\]